OpenStudy (swag):

Does (2x-5) · (x+1) equal 2x^2-5 if not why @Mertsj @campbell_st

OpenStudy (mertsj):

It does not because you forgot the Outer and Inner of FOIL

OpenStudy (campbell_st):

nope... you will have a trinomial... 3 terms you are missing the middle term \[-5 \times x.... and .... 2x \times 1\] these are like terms so can be added or subtracted to find the 3rd term

OpenStudy (campbell_st):

you have \[2x \times x .... and -5 \times 1 \]

OpenStudy (mertsj):